# Synthesizing Strong-Coupling Kohn-Luttinger Superconductivity in 2D Van der Waals materials

### v1

@@ Abstract
Near half-filling, we find strong density fluctuations in DMFT calculations despite no stable long-range ordered state found on the triangular lattice. For comparison, in Fig. 4c, we present a similar phase diagram for the non-frustrated cubic lattice (where the layer index $l$ now can be understood as a pseudo-spin index). At half-filling ($n=0.5$), this system indeed develops a density wave (DW) phase (Fig. 4c, see also Appendix). This long-range DW instability competes with SC order, which can be destroyed by doping electrons (holes). As DW order vanishes (e.g. $n\gtrsim 0.6$ in Fig. 4c), SC instability emerges. Further increasing $n$ leads to the decreasing of superconducting $T_{c}$, similar to the triangular lattice result. This finding demonstrates that strong-coupling KL SC is robust against changes in lattice geometry (Cubic V.S. Triangular) and dimensionality (3D V.S. 2D).
Figure 5: Effective pairing attraction $V^{*}$ as a function long-range Coulomb repulsion $U_{nn}$ for spinless and spinfull models on 2d Triangular lattice at $U=4,U_{o}=8,U^{\prime}=0,T=1.0$. $U_{nn}$ denotes repulsion between intra-layer nearest-neighbouring sites, $U_{nn}\equiv U_{ll}(|r|=1)$.
@@
Spinful model and effects of $U_{lm}(r)$ - In the spinful model of Eq. 1, twofold competing effects can influence interlayer KL pairing relative to the spinless case: (1) the extra spin degree of freedom in the middle layer can enhance pairing by offering an additional polarization channel mediating attraction; (2) on the other hand, strong on-site repulsion $U_{o}$ between opposite-spin electrons induces local electron–hole pairs ($i.e.$, local moments), which is detrimental to interlayer electron-electron pairing. How $T_{c}$ of KL SC changes in the spinful model is therefore determined by the balance between these two opposing mechanisms.
@@
We examine the influence of both the on-site $U_{o}$ and long-range $U_{lm}(r)$ on $V^{*}$ in Fig. 5. For $U_{lm}(r)$, as a concrete example, we consider its intra-layer nearest-neighbor (NN) component, denoted as $U_{nn}\equiv U_{ll}(|r|=1)$. Fig. 5a shows that at $U_{nn}=0$, the spinfull model (at $U_{o}=8t,U=4t$) yields a significantly smaller effective attraction $V^{*}$ compared to the corresponding spinless model ( dashed line, without $U_{o}$, and $U=4t,U_{nn}=0$), confirming that $U_{o}$ suppresses KL pairing. As $U_{nn}$ increases, $V^{*}$ grows in the spinful case, and it exceeds the spinless value once $U_{nn}>2t$. This finding indicates that $U_{nn}$ in fact boosts KL pairing. This is because long-range Coulomb repulsions generally promotes charge density wave (CDW) fluctuations, which intrinsically suppresses the detrimental local moment effect induced by $U_{o}$. At a higher filling $n=0.8$ (Fig. 5b), $V^{*}$ is also seen enhanced by $U_{nn}$. In this case, the spinful model already gives a larger $V^{*}$ than its spinless counterpart even at $U_{nn}=0$. This is expected, because, far away from half-filling ($n\gg 0.5$), local moments are suppressed by doping, thus the additional polarization channel effect in the spinful model becomes predominant, which results in a large $V^{*}$.
@@
Figure 6: Crystal and band structure of the Iron-doped Phosphorene and decorated transition metal halides $\mathrm{CrCl_{3}}$. (a) Top view of the crystal structure of $\mathrm{P_{35}Fe}$. (b) Energy band structure of $\mathrm{P_{35}Fe}$. (c) Crystal structure of $\mathrm{Na}$-absorbed $\mathrm{CrCl_{3}}$, $\mathrm{NaCr_{2}Cl_{6}}$. (d) Energy band structure of $\mathrm{NaCr_{2}Cl_{6}}$.
@@
Realization of strong-coupling KL pairing- As demonstrated above, enhanced KL pairing can be achieved in a trilayer structure of 2D electrons when interlayer repulsion $U$ is comparable to the bandwidth $W$. In 2D van der Waals layered materials, the interlayer interaction can reach $U\lesssim 0.5eV$ (see Appendix and Ref. [63]). To satisfy the strong-coupling condition $U\gtrsim W$, a narrow electronic bandwidth is therefore needed, $W\lesssim 0.5eV$. Through density functional theory (DFT) and constrained random-phase approximation (cRPA) calculations, we have identified several atomically thin two-dimensional half-metals that can effectively realize the trilayer spinless Hubbard model fulfilling $W\lesssim U$. Examples include iron-doped phosphorene ($\mathrm{FeP_{35}}$), with $W\sim 0.37eV$ and $U\sim 0.15eV$ , and sodium-adsorbed transition-metal trihalides such as $\mathrm{NaCr2Cl_{6}}$, where $W\sim 0.087eV$ and $U\sim 0.3eV$ (see Fig. 6 and Appendix). Finally, recent advances in engineering artificial two-dimensional electron lattices in semiconductor heterostructures have achieved bandwidths of only a few meV [65], offering another promising platform for realizing strong KL superconductivity.
@@
Discussion and conclusion - We have proposed that multi-layered 2D electron systems can be engineered to realize a novel emergent quantum many-body state: strong-coupling KL superconductivity. A key advantage of our proposal lies in the large pairing energy scale of the strong-coupling KL SC, which does not need an intermediate collective mode [64, 75] to mediate the pairing. As a result, it exhibits robustness against variations in lattice geometry and dimensionality, and can endure a large remnant Coulomb repulsion $U^{*}$ between pairing electrons. These advantages allow substantial flexibility to control the pairing by tuning sample thickness, dielectric environment or interlayer distance in VdW materials. Although our low-temperature DMFT calculations of $T_{c}$ are restricted to the spinless model for half-metals due to numerical challenges, the DQMC results do suggest that the spinful model (for normal 2D metals) can, under proper conditions, host an even stronger effective pairing attraction $V^{*}$.
@@
Acknowledgment We are grateful to Fadi Sun for useful discussions. This work is supported by the National Natural Science Foundation of China (Grants No.12274472, No. 12494594). We also thank the support from the Research Center for Magnetoelectric Physics of Guangdong Province (Grants No. 2024B0303390001) and Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2401010).

### v2

@@ Abstract
In vicinity of half-filling, we find strong density fluctuation in CDMFT calculations which tends to form staggered distribution of electrons at the three layers ( see Fig. 5b ). However, due to the geometry frustration, no true long-range ordered state is found on triangular lattice. For comparison, in Fig. 4c, we present a similar phase diagram for the non-frustrated cubic lattice (where the layer index $l$ now can be understood as a pseudo-spin index). At half-filling ($n=0.5$), this system indeed develops a density wave (DW) phase. This long-range DW instability competes with SC order, which can be destroyed by doping electrons (holes). As DW order vanishes (e.g. $n\gtrsim 0.6$ in Fig. 4c), SC instability emerges. Further increasing $n$ leads to the decreasing of superconducting $T_{c}$, similar to the triangular lattice result. This finding demonstrates that strong-coupling KL SC is robust against changes in lattice geometry (bipartite V.S. non-bipartite) and dimensionality (3D V.S. 2D).
Figure 5: Long-range Coulomb repulsions can enhance pairing attraction $V^{*}$. $U_{nn}=U_{ll}(\vec{r_{1}})$ denotes the intra-layer NN repulsion, $\vec{r_{1}}||_{xy}=1$.(a) $V^{*}$ as a function of $U_{nn}$ for spinless model at $T=1.0$. Other NN repulsions are set as $U_{12}(\vec{r_{1}})=U_{23}(\vec{r_{1}})=U_{nn}/2$, $U_{13}(\vec{r_{1}})=U_{nn}/4$. (b) A staggered distribution of electrons at three layers minimizing the potential energy is shown in 1D chain, on which interaction terms that promote (solid lines) or suppress (dashed lines) inter-layer pairing is illustrated. Filled (hollow) symbols denote electrons (holes). (c) and (d): $V^{*}$ as a function of $U_{nn}$ for the spinful model at $U_{o}=8,U=4,U^{*}=0,T=1.0$. Inter-layer NN repulsion $U_{lm}(\vec{r_{1}}),l\neq m$ is neglected in (c) and (d) to avoid sign problem.
Fffects of $U(r)$ and spinful model- Considering long-ranged Coulomb repulsions beyond a unit cell, there can be different $U_{lm}(r)$ terms stabilizing or destabilizing the binding of electron (hole) pairs at top- and bottom- layers, as illustrated in Fig. 5b. We find that the overall effects of the nearest-neighboring (NN) repulsion enhances $V^{*}$, as shown in Fig. 5a for the spinless model.
@@
For the spinful model of Eq. 1, twofold competing effects can influence interlayer KL pairing relative to the spinless case: (1) the extra spin degree of freedom in the middle layer can enhance pairing by offering an additional polarization channel; (2) on the other hand, strong on-site repulsion $U_{o}$ between opposite-spin electrons induces local electron–hole pairs ($i.e.$, local moments), which is detrimental to interlayer electron-electron pairing. The balance between these two opposing mechanisms then affects $T_{c}$ in the spinful model.
@@
We examine influence of both on-site $U_{o}$ and NN repulsion $U_{nn}$ on $V^{*}$ in Fig. 5c and Fig. 5d at $U_{o}=8t,U=4t$. Fig. 5c shows that at $U_{nn}=0$, the spinfull model yields a smaller $V^{*}$ compared to the corresponding spinless model ( dashed line, without $U_{o}$, and $U=4t,U_{nn}=0$), confirming that $U_{o}$ suppresses KL pairing. As $U_{nn}$ increases, $V^{*}$ grows and exceeds the spinless value once $U_{nn}\gtrsim 2t$. We note that long-range Coulomb repulsions can promote charge density wave (CDW) fluctuations, hence suppressing the detrimental local moments induced by $U_{o}$. This could be another route that $U(r)$ enhances $V^{*}$ in spinful cases. At a higher filling $n=0.8$ (Fig. 5b), $V^{*}$ is also seen enhanced by $U_{nn}$. In this case, the spinful model already gives a larger $V^{*}$ than its spinless counterpart even at $U_{nn}=0$. This is expected, because, far away from half-filling, local moments are suppressed by doping, thus the additional polarization channel effect in spinful model becomes predominant, which results in a larger $V^{*}$.
@@
Figure 6: Crystal and band structure of $\mathrm{Na}$-adsorbed transition metal halides $\mathrm{CrCl_{3}}$. (a) Crystal structure of $\mathrm{Na}$-adsorbed $\mathrm{NaCr_{2}Cl_{6}}$. (b) Energy band structure of $\mathrm{NaCr_{2}Cl_{6}}$. Inset: Magnified view of energy bands near Fermi level.
@@
Realization of strong-coupling KL pairing- As demonstrated above, enhanced KL pairing can be achieved in a trilayer structure of 2D electrons when $U\sim W$. In 2D van der Waals layered materials, the interlayer repulsion can reach $U\lesssim 0.5eV$ (see End Matter and Ref. [63]). To satisfy the strong-coupling condition, a narrow electronic bandwidth, $W\lesssim 0.5eV$, is therefore needed. Through density functional theory (DFT) and constrained random-phase approximation (cRPA) calculations, we have identified several atomically thin two-dimensional half-metals that can effectively realize the trilayer spinless Hubbard model fulfilling $W\lesssim U$. Examples include sodium-adsorbed transition-metal trihalides such as $\mathrm{NaCr_{2}{C}l_{6}}$, where $W\sim 0.087eV$ and $U\sim 0.3eV$ (see Fig. 6 and End Matter), and an iron-doped phosphorene, $\mathrm{FeP_{35}}$ [34] (see End Matter). We note that the transition metal dichalcogenide(TMD)-based moiré superlattices [78] exhibit bandwidths below $10\mathrm{meV}$ (inter-site hopping strength $t\sim 1\mathrm{meV}$[67], which is smaller than the inter-layer repulsion [72]. Hence these systems may also be suitable for implementing our proposal. Finally, recent advances in engineering artificial 2D electron lattices within semiconductor heterostructures have yielded bandwidths of only a few meV [65], offering another promising platform for realizing strong KL superconductivity.
@@
Discussion and conclusion - We have proposed a scheme for realizing a novel emergent quantum many-body state: strong-coupling KL superconductivity in 2D electrons. Recently, novel exciton-mediated SC in atomically thin VdW heterostructures have been proposed [64, 77]. A key distinction of our proposal lies in the large pairing energy scale $V^{*}$ that does not require any collective mode to mediate the pairing. Consequently, it exhibits robustness against variations in lattice geometry and dimensionality, and can endure a large remnant Coulomb repulsion $U^{*}$ between pairing electrons. These advantages allow substantial flexibility to control SC by tuning sample thickness, dielectric environment or interlayer distance in VdW materials. Although our low-temperature DMFT calculations of $T_{c}$ are restricted to the spinless model for half-metals due to numerical challenges, the DQMC results do suggest that the spinful model (for normal 2D metals) can, under proper conditions, host an even stronger effective pairing attraction $V^{*}$.
@@
Acknowledgment We are grateful to Wenchen Luo and Fadi Sun for useful discussions. This work is supported by the National Natural Science Foundation of China (Grants No.12274472, No. 12494594). We also thank the support from the Research Center for Magnetoelectric Physics of Guangdong Province (Grants No. 2024B0303390001) and Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2401010). H.Y. acknowledges support by Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2501003) and NSFC under Grant No. 12274477.